Consider a bond whose annual coupon rate is 10% and coupons are paid twice a year evenly. Its face-value is $100,000 and maturity is 2 years. Yield-to-maturity is 10% (annual) (fixed). What are the duration (in years) and convexity of the bond?
Answer #1
                  K = Nx2
Bond Price =∑ [( Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =2x2
Bond Price =∑ [(10*100000/200)/(1 + 10/200)^k]     +   100000/(1 + 10/200)^2x2
                   k=1
Bond Price = 100000

php0EPQZ3.png

C =.

Period Cash Flow Discounting factor PV Cash Flow Duration Calc Convexity Calc
0 ($100,000.00) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period =duration calc*(1+period)/(1+YTM/N)^2
1           5,000.00                                                             1.05              4,761.90              4,761.90              8,638.38
2           5,000.00                                                             1.10              4,535.15              9,070.29            24,681.07
3           5,000.00                                                             1.16              4,319.19            12,957.56            47,011.57
4      105,000.00                                                             1.22            86,383.76          345,535.04        1,567,052.33
      Total          372,324.80        1,647,383.35
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=372324.8/(100000*2)
=1.861624
Modified duration = Macaulay duration/(1+YTM)
=1.86/(1+0.1)
=1.772975
Convexity =(∑ convexity calc)/(bond price*number of coupon per year^2)
=1647383.35/(100000*2^2)
=4.12